OpenStudy (anonymous):

y = (3/8)x - 7 (9, -4) (8, -4) (7, -4) (8, -5)

OpenStudy (anonymous):

(8, -4) you get it from plugging in the points and determining if they are equal. for instance if you plug in 9 for x and -4 for y you will get 27/8 - 7 = -4 which will never happen

OpenStudy (anonymous):

Determine the slope of the line that passes through the points (-1, 3) and (1, 5).

OpenStudy (anonymous):

use the formula \[m = (y2 - y1) / (x2 - x1) \] and then plug back in a point to get the y-intercept b and then you write the general formula y = mx + b

OpenStudy (anonymous):

thanks!